# Multilingual large language models leak human stereotypes across language boundaries

### v1

@@ 3.2.2. Negative Leakage
On the other hand, there are negative stereotypes that leak across languages. From the results of the English and Russian survey, “feminists” are viewed more as cold, while in Hindi they are perceived more as warm. We observe a leakage from Russian and English languages to Hindi enforcing negative stereotypes about feminists. Another example is “civil servants”. Historically, this social group is more negatively viewed in Russia, China, and India compared to the United States. People related to the government are typically viewed as wealthy and dominant, which we observe to leak to English. Simultaneously we observe a positive leakage from English to Russian and Chinese for the trait likable, which also confirms different perceptions between annotators for these languages. There is a notable leakage from English to Russian, Chinese, and Hindi for “Black people” for such traits as religious, unassertive, dominated, low status. This aligns with known stereotypes about African Americans and Africans in U.S. society (Miller-Cribbs and Farber, 2008; Galster, 1992; BERESFORD, 1996).
@@ 3.2.3. Non-polar Leakage
There are also non-polar leakages, which are neither positive nor negative. From Chinese to Hindi and English, we see the strengthening of non-religious trait for various groups such as “software engineers”, “veterans”, “wealthy people”, and “government officials”. It has been shown that there are non-believers of the total population in China as of (Yang and Huang, 2018).
@@ 3.2.4. Non-shared Groups Leakage
In the case of non-shared groups, we expected uni-directional transferring of the groups’ perceptions from the language of origin to other languages. Our findings confirm this hypothesis. For example, the group “VDV soldiers” is a widely known military unit in Russia. There are strong stereotypes in Russian society about this group, but the group is mostly unknown to Americans. Out of the survey English survey respondents who passed the quality tests, no one chose this group as a familiar one. This group’s representation leaks from Russian to English, strengthening traits such as trustworthy, sincere, threatening, nonreligious, and confident. Another example is “Hui people”, a group widely unknown to Russian and Hindi society: out of respondents for both surveys, no one chose this group as the familiar one. This social group is a minority in China and is composed of Chinese-speaking followers of Islam. Originally, “Hui people” are marginalized in China and viewed as more traditional, religious, and conservative (Hillman, 2004; Hong, 2005). Accordingly, we observed the leakage of such traits as conservative, traditional, religious, egoistic. All groups specific to the Hindi language — “Gujarati, Brahmin”, and “Shudra people” — have certain traits leaking to the English language. For example, high caste groups (“Gujarati” and “Brahmin people”) strengthen such positive traits as wealthy, powerful, high status, likable, sincere, trustworthy. In addition, “Brahmin people” become more associated in ChatGPT with traits traditional, dominant. “Shudra people” become more associated with the trait unassertive. This leakage corresponds to the perception of these groups in Indian society and by our survey respondents (Witzel, 1993; Milner, 1993).
@@ 4. Conclusion & Limitations
Multilingual large language models have the potential to spread stereotypes beyond the societal context they emerge from, whether by generating new stereotypes, amplifying existing ones, or reinforcing prevailing social perceptions from dominant cultures. In our work, we demonstrate that this concern is indeed valid. To do so, we establish a framework for measuring the leakage of stereotypical associations in multilingual large language models across languages. We are limited in our ability to run a causal analysis, because none of the studied languages can be easily removed from the training data to see their genuine impact on stereotypical associations in other languages. Retraining ChatGPT, for instance, is not a feasible option. Nonetheless, if our hypothesis holds, indicating that stereotype leakage is occurring, we would anticipate observing associations of stereotypes cross-lingually, and indeed, we do identify this association. As a proxy for re-training the models without a particular language, we run an association on a monolingual version of the model. We perform this experiment on both monolingual BERTs and multilingual BERT to measure how well the monolingual BERT in a specific language can predict the behavior of mBERT in the same language. We find stronger associations between stereotypes in monolingual English and Russian models with mBERT in the same languages than for the case of Chinese and Hindi languages. In addition, we find that there is interaction and exchange between languages in multilingual large language models. The stereotype leakage occurs bidirectionally. On the example of ChatGPT, as the best-aligned with human judgment model, we observe the strengthening of positive, negative, and non-polar associations in the model. In addition, our study underscores the role of “native” languages in framing social groups unknown to other linguistic communities. Such leakage of stereotypes amplifies the complexity of societal perceptions by introducing a complex interconnected bias from different languages and cultures. In the context of shared groups, stereotype leakage may manifest as the manifestation of stereotypes that were not previously present within the cultural setting of a particular group. In the case of non-shared groups, stereotype leakage can extend the reach of existing stereotypes from the source culture to other cultural contexts.
@@
Our work has several limitations. First, stereotypes were selected based on the ABC model, which was developed and tested using U.S. and German stereotypes. We translated our surveys into the other languages but this might result in patterns that better reflect Anglocentric stereotypes (Talat et al., 2022) than other stereotypes. In our study, we try to control the influence of the U.S. culture by asking crowd workers how frequently they read U.S. social media. We see that on average of respondents in Russia, China, and India read the media. This American cultural dominance might affect the collected data as these data may not fully capture the range of stereotypes typical for these cultures. In addition, we have language-culture limitations as English language survey results only apply to the U.S., Russian to Russia, Chinese to China, and Hindi to India. Lastly, while we indirectly consider culture through survey results on associations, we do not measure or account for culture in a comprehensive manner.
@@ Acknowledgements.
- BERESFORD (1996) PETER BERESFORD. 1996. Poverty and Disabled People: Challenging dominant debates and policies. Disability & Society 11, 4 (1996), 553–568. https://doi.org/10.1080/09687599627598 arXiv:https://doi.org/10.1080/09687599627598
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, and Dario Amodei. 2020. Language Models are Few-Shot Learners.
@@
- Câmara et al. (2022) António Câmara, Nina Taneja, Tamjeed Azad, Emily Allaway, and Richard Zemel. 2022. Mapping the Multilingual Margins: Intersectional Biases of Sentiment Analysis Systems in English, Spanish, and Arabic. arXiv:2204.03558 [cs.CL]
- Galster (1992) George Galster. 1992. Housing Discrimination and Urban Poverty of African-Americans. Journal of Housing Research 2 (01 1992).
@@
- Hillman (2004) Ben Hillman. 2004. The rise of the community in rural China: Village politics, cultural identity and religious revival in a Hui hamlet. The China Journal 51 (2004), 53–73.
@@
- Ouyang et al. (2022) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L. Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simens, Amanda Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe. 2022. Training language models to follow instructions with human feedback. arXiv:2203.02155 [cs.CL]
- Pieke (2012) Frank N. Pieke. 2012. Immigrant China. Modern China 38, 1 (2012), 40–77. https://doi.org/10.1177/0097700411424564
@@
- THELWELL (2020) CHINUA THELWELL. 2020. Front Matter. University of Massachusetts Press, i–iv. http://www.jstor.org/stable/j.ctv160btb3.1
@@
- Yang and Huang (2018) Yu Yang and Shizhi Huang. 2018. Religious Beliefs and Environmental Behaviors in China. Religions 9, 3 (2018). https://doi.org/10.3390/rel9030072

### v2

@@ 3.2.2. Negative Leakage
On the other hand, there are negative stereotypes that leak across languages. From “feminists”, we observe a leakage from English to Chinese and Hindi and from Russian to Chinese of such stereotypical associations as egoistic, threatening, repellent, and cold, while, for instance, in the human data in Hindi this group is perceived as warm.
Another example is “immigrants”. From Russian and English languages, traits such as threatening, repellent, dishonest, egoistic, and unconfident leak to Chinese and Hindi. Based on human data, we found that people surveyed in Chinese view this group quite favorably since the majority of immigrants to China are highly qualified professionals (Pieke, 2012). In Russia, immigrants are mostly coming from poorer neighboring countries and are negatively stereotyped in society, while in the U.S., immigrants are diverse and could be both marginalized or privileged.
@@ 3.2.3. Non-polar Leakage
There are also non-polar leakages, which are neither positive nor negative. From Hindi to English and Russian, we see the strengthening of religious for various groups such as “women”, “disabled people”, “Black people”, and “Asian people”. It has been shown that there are more than believers of the total population in India as of (Sahgal et al., 2021).
@@ 3.2.4. Non-shared Groups Leakage
In the case of non-shared groups, we expected uni-directional transferring of the groups’ perceptions from the language of origin to other languages. Our findings confirm this hypothesis. For example, the group “VDV soldiers” is a widely known military unit in Russia. There are strong stereotypes in Russian society about this group, but the group is mostly unknown to Americans. Out of the survey English survey respondents who passed the quality tests, no one chose this group as a familiar one. This group’s representation leaks from Russian to English, strengthening traits such as confident, traditional, competitive, and threatening. Another example is “Hui people”, a group widely unknown to Russian and Hindi society: out of respondents for both surveys, no one chose this group as the familiar one. This social group is a minority in China and is composed of Chinese-speaking followers of Islam. Originally, “Hui people” are marginalized in China and viewed as more traditional, religious, and conservative (Hillman, 2004; Hong, 2005). Accordingly, we observed the leakage of such traits as irrational, traditional, threatening, repellent, religious, and egoistic. All groups specific to the Hindi language — “Gujarati, Brahmin”, and “Shudra people” — have certain traits leaking to the English and Russian languages. For example, high caste groups (“Gujarati” and “Brahmin people”) strengthen such positive traits as wealthy, likable, sincere, powerful, high status, competitive, and confident. In addition, “Brahmin people” become more associated in GPT-3.5 with traits poor, low status, powerless, traditional, religious, and dominated. This leakage corresponds to the perception of these groups in Indian society and by our survey respondents (Witzel, 1993; Milner, 1993).
3.2.5. Discussion
The amplification of negative stereotypes is certainly a cause for concern. These stereotypes, often deeply ingrained in societal narratives, can perpetuate discrimination and prejudice. Conversely, while positive stereotypes might seem harmless or even beneficial at first glance, they can also be problematic. In some contexts, positive stereotypes may serve to counterbalance negative ones, creating what is known as an anti-stereotype effect. This can be useful in mitigating some of the harms caused by negative stereotypes.
However, positive stereotypes, such as the notion that “Asian people” are wealthy or “housewives” are warm, can also lead to unrealistic expectations and pressures. For instance, not all Asian people are wealthy, and assuming so can ignore the diverse economic realities faced by individuals within this broad demographic. Similarly, the stereotype that housewives are inherently warm can enforce restrictive roles based on gender.
The leakage of stereotypes is particularly troubling for certain applications, including education and creative content generation. These fields heavily influence public perception and personal development, making the integrity of the content they deliver crucial. Systems built for these applications with MLLMs must be particularly cautious of the stereotype leakage effect. It is essential for developers to implement strategies that actively mitigate the harmful leakage effects.
@@ 4. Conclusion & Limitations
Multilingual large language models have the potential to spread stereotypes beyond the societal context they emerge from, whether by generating new stereotypes, amplifying existing ones, or reinforcing prevailing social perceptions from dominant cultures. In our work, we demonstrate that this concern is indeed valid. To do so, we establish a framework for measuring the leakage of stereotypical associations in multilingual large language models across languages. Overall, we find that the stereotype leakage occurs bidirectionally meaning that when one language transmits stereotypes to others, it likely receives some stereotypes from other languages as well. We also observe the most stereotype leakage effect within the GPT-3.5 model. Within the GPT-3.5 model, we observe the strengthening of positive, negative, and non-polar associations in the model. In addition, our study underscores the role of “native” languages in framing social groups unknown to other linguistic communities. Such leakage of stereotypes amplifies the complexity of societal perceptions by introducing a complex interconnected bias from different languages and cultures. In the context of shared groups, stereotype leakage may manifest as the manifestation of stereotypes that were not previously present within the cultural setting of a particular group. In the case of non-shared groups, stereotype leakage can extend the reach of existing stereotypes from the source culture to other cultural contexts.
@@
Our work has several limitations. First, we are limited in our ability to run a causal analysis because none of the studied languages can be easily removed from the training data to see their genuine impact on stereotypical associations in other languages. Retraining GPT-3.5, for instance, is not a feasible option. Thus, we use BERT monolingual model as a proxy for each language. In addition, stereotype traits were selected based on the ABC model, which was developed and tested using U.S. and German stereotypes. Though we translated our surveys into all four languages, the stereotype traits may better reflect Anglocentric stereotypes (Talat et al., 2022) than others. Furthermore, the human stereotypes we collected may already reflect the influence of social stereotype transmission. For instance, in our study, we surveyed crowd workers about their consumption of U.S. social media. We found that, on average, of respondents from Russia, China, and India engage with U.S. social platforms. Such American cultural dominance could potentially affect the human stereotypes collected in these three languages. Lastly, while we indirectly consider culture through survey results on associations, we do not measure or account for culture in a comprehensive manner. Our English language survey results only apply to the U.S., Russian to Russia, Chinese to China, and Hindi to India. Lastly,
@@ Acknowledgements.
- Beresford (1996) Peter Beresford. 1996. Poverty and Disabled People: Challenging dominant debates and policies. Disability & Society 11, 4 (1996), 553–568. https://doi.org/10.1080/09687599627598
- Brown et al. (2020) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language Models are Few-Shot Learners. CoRR abs/2005.14165 (2020). arXiv:2005.14165 https://arxiv.org/abs/2005.14165
@@
- Câmara et al. (2022) António Câmara, Nina Taneja, Tamjeed Azad, Emily Allaway, and Richard Zemel. 2022. Mapping the Multilingual Margins: Intersectional Biases of Sentiment Analysis Systems in English, Spanish, and Arabic. In Proceedings of the Second Workshop on Language Technology for Equality, Diversity and Inclusion, Bharathi Raja Chakravarthi, B Bharathi, John P McCrae, Manel Zarrouk, Kalika Bali, and Paul Buitelaar (Eds.). Association for Computational Linguistics, Dublin, Ireland, 90–106. https://doi.org/10.18653/v1/2022.ltedi-1.11
@@
- Galster and Carr (1991) George C. Galster and James H. Carr. 1991. Housing Discrimination and Urban Poverty of African-Americans. Journal of Housing Research 2, 2 (1991), 87–123. http://www.jstor.org/stable/24825920
@@
- Hillman (2004) Ben Hillman. 2004. The Rise of the Community in Rural China: Village Politics, Cultural Identity and Religious Revival in a Hui Hamlet. The China Journal 51 (2004), 53–73. http://www.jstor.org/stable/3182146
@@
- Pieke (2012) Frank N. Pieke. 2012. Immigrant China. Modern China 38, 1 (2012), 40–77. http://www.jstor.org/stable/23216934
@@
- Sahgal et al. (2021) Neha Sahgal, Jonathan Evans, Ariana Monique Salazar, Kelsey Jo Starr, and Manolo Corichi. 2021. Religion in India: Tolerance and Segragation. Pew Research Centre (2021). https://doi.org/202.419.4372
@@
- Thelwell (2020) Chinua Thelwell. 2020. Exporting Jim Crow: Blackface Minstrelsy in South Africa and Beyond. University of Massachusetts Press. http://www.jstor.org/stable/j.ctv160btb3